### v1

@@ Abstract
This paper investigates a Halpern acceleration of the inexact proximal point method for solving maximal monotone inclusion problems in Hilbert spaces. The proposed Halpern inexact proximal point method (HiPPM) is shown to be globally convergent, and a unified framework is developed to analyze its worst-case convergence rate. Under mild summability conditions on the inexactness tolerances, HiPPM achieves an $\mathcal{O}(1/k^{2})$ rate in terms of the squared fixed-point residual. Furthermore, under additional mild condition, the method retains a fast linear convergence rate. Building upon this framework, we further extend the acceleration technique to constrained convex optimization through the augmented Lagrangian formulation. In analogy to Rockafellar’s classical results, the resulting accelerated inexact augmented Lagrangian method inherits the convergence rate and complexity guarantees of HiPPM. The analysis thus provides a unified theoretical foundation for accelerated inexact proximal algorithms and their augmented Lagrangian extensions.
@@ 1 Introduction
where $T:\mathcal{H}\rightrightarrows\mathcal{H}$ is a maximally monotone operator. This problem provides a unifying framework for a broad range of models in optimization and equilibrium theory. In particular, many convex optimization problems, variational inequalities, and saddle-point formulations can be recast as finding the zeros of a monotone operator. The proximal point method (PPM) is a fundamental approach for solving monotone inclusion problems of the form (1.1). It not only enjoys strong global convergence guarantees for maximal monotone operators, but also serves as the theoretical foundation for various operator-splitting schemes. In particular, for convex optimization problems, both the augmented Lagrangian method (ALM) and the alternating direction method of multipliers (ADMM) can be interpreted as instances of the PPA when applied to their respective dual formulations.
@@
$P_{c}(z)=(I+cT)^{-1}(z),z\in{\cal H}.$ (1.2)
It follows from [10, Proposition 1] that $P_{c}$ is a nonexpansive mapping with
$0\in T(z)\mbox{ if and only if }z=P_{c}(z),$
then or equivalently,
@@
In contrast to Nesterov-type acceleration, which is originally developed for convex optimization, Halpern’s fixed-point iteration [4] is designed for approximating fixed points of nonexpansive operators. Since Contreras and Cominetti [2] revealed a close connection between Kim’s accelerated proximal point method [5] and Halpern’s iteration, it is natural to consider Halpern-type schemes for accelerating the proximal point method, given their simplicity. Lieder [6] studied Halpern’s iteration in Hilbert spaces and establish the $\mathcal{O}(1/k^{2})$ convergence rate. Moreover, under the co-coercive assumption, Tran-Dinh [11] derive an equivalence between the Nesterov’s accelerate method and Halpern’s fixed-point iteration. However, these connections do not directly extend to settings in which the subproblems are solved inexactly, which is typically the case in practice.
In this paper, we study a Halpern-type acceleration of the inexact proximal point method (PPM) for solving the maximal monotone inclusion problem (1.1) in Hilbert spaces. The main contributions of this work can be summarized as follows:
@@
    As pointed out in [5], for strongly monotone operators, the standard proximal point method enjoys a linear convergence rate in terms of the fixed-point residual, whereas its Halpern-accelerated variant is not necessarily guaranteed to preserve such a rate. In this paper, we demonstrate that the inexact PPM with Halpern acceleration indeed attains linear convergence under the strong monotonicity assumption.
@@
The remainder of this paper is organized as follows. Section 2 introduces the framework of the Halpern accelerated inexact proximal point algorithm ( in Hilbert spaces and establishes its global convergence. Section 3 develops a general framework for analyzing the worst-case convergence rate of the proposed HiPPA using a direct algebraic proof, and demonstrates its fast linear convergence under mild conditions. Section 4 further derives the convergence rate of the Halpern accelerated inexact augmented Lagrangian method, based on its close connection with the Halpern accelerated inexact PPA.
@@
2 Halpern acceleration of the inexact PPA
Building upon Rockafellar’s inexact proximal point algorithm [9], we introduce an additional Halpern-type acceleration step to enhance its convergence behavior. The resulting accelerated scheme, referred to as HiPPA, can be formulated as follows.
@@ 2 Halpern acceleration of the inexact PPA
We recall the following result from [10, Proposition 2], which serves as a fundamental criterion for the existence of solutions to the inclusion problem $0\in T(z)$.
@@
For notational convenience,, for $k=0,1,\ldots$, define
$\eta^{k}=\bar{z}^{k}-P_{c_{k}}(z^{k}).$ (2.3)
@@
$\|\bar{z}^{k}\|=\left\|\displaystyle\frac{1+2}{k+1}z^{k+1}-\displaystyle\frac{1}{k+1}z^{0}\right\|\leq 1.4s+\left\|\displaystyle\frac{1}{k+1}z^{0}\right\|<1.5s+0.5s=2s,$
@@
Then, the result follows by applying the same technique as in [10, Theorem 1]. For completeness, the details are provided below. Consider the set-valued mapping $T^{\prime}$ defined by
@@
Thus the sequence $\{z^{k}\}$ can be regarded as arising from the Halpern accelerated proximal point algorithm for operator $T^{\prime}$. As the effective domian ${\rm Dom}\,(T^{\prime})$ is bounded, so that $(T^{\prime})^{-1}(0)\neq\emptyset$ by Lemma 2.1. It follows from (2.5) that $(T^{\prime})^{-1}(0)\subseteq T^{-1}(0)$, this implies $T^{-1}(0)\neq\emptyset$. The proof is completed. ∎
@@
$\lim_{k\rightarrow\infty}\|Q_{c_{k}}(z^{k})\|=0,\quad\lim_{k\rightarrow\infty}\|\bar{z}^{k}-z^{k}\|=0.$ (2.8)
@@
$\|z^{k}\|<s\mbox{ and }\varepsilon_{k}<s,\quad\forall k.$ (2.9)
@@
$\|\hat{z}^{k}\|\leq 2s,\quad\|P_{c_{k}}(z^{k})\|\leq 3s,\quad\mbox{ when }k>N_{0}.$
@@
$\begin{array}[]{ll}\|Q_{c_{k}}(z^{k})\|^{2}-\|z^{k}-z^{0}\|^{2}+\|z^{k+1}-z^{0}\|^{2}&\leq\|z^{k+1}-z^{0}\|^{2}-\|P_{c_{k}}(z^{k})-z^{0}\|^{2}\\[8.0pt] &=\left(\displaystyle\frac{k+1}{k+2}\right)^{2}\|z^{0}-\bar{z}^{k}\|^{2}-\|(P_{c_{k}}(z^{k})-\bar{z}^{k})+(\bar{z}^{k}-z^{0})\|^{2}\\[8.0pt] &\leq-\|(P_{c_{k}}(z^{k})-\bar{z}^{k})\|^{2}-2\langle(P_{c_{k}}(z^{k})-\bar{z}^{k}),(\bar{z}^{k}-z^{0})\rangle\\[8.0pt] &\leq 2\|P_{c_{k}}(z^{k})-\bar{z}^{k}\|\|\bar{z}^{k}-z^{0}\|\leq 2(\|z^{0}\|+2s)\varepsilon_{k}.\end{array}$
@@
$\begin{array}[]{ll}\|Q_{c_{k}}(z^{k})\|^{2}-\|z^{k}-\bar{z}\|^{2}+\|z^{k+1}-\bar{z}\|^{2}&\leq\|z^{k+1}-\bar{z}\|^{2}-\|P_{c_{k}}(z^{k})-\bar{z}\|^{2}\\[8.0pt] &=\left\|\displaystyle\frac{1}{k+2}(z^{0}-\bar{z})+\displaystyle\frac{k+1}{k+2}(\bar{z}^{k}-\bar{z})\right\|^{2}-\|P_{c_{k}}(z^{k})-\bar{z}^{k}+\bar{z}^{k}-\bar{z}\|^{2}\\[8.0pt] &\leq\displaystyle\frac{1}{(k+2)^{2}}\|(z^{0}-\bar{z})\|^{2}-2\langle P_{c_{k}}(z^{k})-\bar{z}^{k},\bar{z}^{k}-\bar{z}\rangle-\|P_{c_{k}}(z^{k})-\bar{z}^{k}\|^{2}\\[8.0pt] &\leq\displaystyle\frac{1}{(k+2)^{2}}\|(z^{0}-\bar{z})\|^{2}+2(2s+\|\bar{z}\|)\varepsilon_{k}.\end{array}$
@@
Arguing as in the proof of [10, Theorem 1], one can show that the sequence $\{z^{k}\}$ admits at most one weak cluster point. Hence, the entire sequence $\{z^{k}\}$ converges weakly to some $z^{\infty}$. ∎
@@ 3 The rate of convergence
$\begin{array}[]{ll}\Delta_{k}&=\displaystyle\frac{4k}{(k+1)^{2}}\left\langle z^{k}-z^{0},\eta^{k-1}\right\rangle+\displaystyle\frac{4}{(k+1)^{2}}\displaystyle\sum_{j=0}^{k-1}\left\langle z^{0}-z^{j},\frac{j+1}{j+2}\eta^{j}\right\rangle\\[6.0pt] &\quad+\displaystyle\frac{4}{(k+1)^{2}}\displaystyle\sum_{j=0}^{k-1}j\left\langle z^{0}-z^{j},\frac{j+1}{j+2}\eta^{j}-\eta^{j-1}\right\rangle\\[6.0pt] &\quad+\displaystyle\frac{4k}{k+1}\left\langle z^{k}-P_{c}(z^{k}),\eta^{k-1}\right\rangle\\[6.0pt] &\quad-\displaystyle\frac{4}{(k+1)^{2}}\displaystyle\sum_{j=0}^{k-1}\displaystyle\frac{(j+1)^{3}}{j+2}\left\langle P_{c}(z^{j})-z^{j},\eta^{j-1}-\eta^{j}\right\rangle\\[6.0pt] &\quad+\displaystyle\frac{4}{(k+1)^{2}}\displaystyle\sum_{j=0}^{k-1}\displaystyle\frac{j+1}{j+2}\left\langle P_{c}(z^{j})-z^{j},\eta^{j-1}\right\rangle\\[6.0pt] &\quad-\displaystyle\frac{4}{(k+1)^{2}}\displaystyle\sum_{j=1}^{k}\displaystyle\frac{j^{2}}{j+1}\|\eta^{j-1}\|^{2}.\end{array}$ (3.2)
@@
$\begin{array}[]{l}j(j+1)\|P_{c}(z^{j})-P_{c}(z^{j-1})\|^{2}\\[4.0pt] =j(j+1)\left\|z^{j}-P_{c}(z^{j})+\displaystyle\frac{1}{j}[z^{j}-z^{0}]-\eta^{j-1}\right\|^{2}\\[4.0pt] =j(j+1)\left[\left\|z^{j}-P_{c}(z^{j})+\displaystyle\frac{1}{j}[z^{j}-z^{0}]\right\|^{2}+\|\eta^{j-1}\|^{2}-2\langle z^{j}-P_{c}(z^{j})+\displaystyle\frac{1}{j}[z^{j}-z^{0}],\eta^{j-1}\rangle\right]\\[4.0pt] =j(j+1)\|z^{j}-P_{c}(z^{j})\|^{2}+2(j+1)\langle z^{j}-P_{c}(z^{j}),z^{j}-z^{0}\rangle+\displaystyle\frac{j+1}{j}\|z^{j}-z^{0}\|^{2}\\[8.0pt] \quad-2j(j+1)\langle z^{j}-P_{c}(z^{j})+\displaystyle\frac{1}{j}[z^{j}-z^{0}],\eta^{j-1}\rangle+j(j+1)\|\eta^{j-1}\|^{2}\\[6.0pt] =A_{1}(j)+A_{2}(j)+A_{3}(j)-2j(j+1)\langle z^{j}-P_{c}(z^{j})+\displaystyle\frac{1}{j}[z^{j}-z^{0}],\eta^{j-1}\rangle+j(j+1)\|\eta^{j-1}\|^{2},\end{array}$ (3.7)
@@
$\begin{array}[]{l}A_{1}(j)=j(j+1)\|z^{j}-P_{c}(z^{j})\|^{2},\,\,A_{2}(j)=2(j+1)\langle z^{j}-P_{c}(z^{j}),z^{j}-z^{0}\rangle,\,\,\hbox{and}\,\,A_{3}(j)=\displaystyle\frac{j+1}{j}\|z^{j}-z^{0}\|^{2}.\end{array}$
@@
$\begin{array}[]{l}-j(j+1)\|z^{j}-z^{j-1}\|^{2}\\[4.0pt] =-j(j+1)\left\|\displaystyle\frac{1}{j+1}(z^{0}-P_{c}(z^{j-1}))+P_{c}(z^{j-1})+\displaystyle\frac{j}{j+1}\eta^{j-1}-z^{j-1}\right\|^{2}\\[8.0pt] =-2j\langle z^{0}-P_{c}(z^{j-1}),P_{c}(z^{j-1})-z^{j-1}\rangle-j(j+1)\|P_{c}(z^{j-1})-z^{j-1}\|^{2}\\[8.0pt] \quad-\displaystyle\frac{j}{j+1}\|z^{0}-P_{c}(z^{j-1})\|^{2}-2j(j+1)\left\langle\displaystyle\frac{1}{j+1}(z^{0}-P_{c}(z^{j-1}))+P_{c}(z^{j-1})-z^{j-1},\displaystyle\frac{j}{j+1}\eta^{j-1}\right\rangle\\[8.0pt] \quad-j(j+1)\left(\displaystyle\frac{j}{j+1}\right)^{2}\|\eta^{j-1}\|^{2}\\[8.0pt] =B_{1}(j)+B_{2}(j)+B_{3}(j)-j(j+1)\left(\displaystyle\frac{j}{j+1}\right)^{2}\|\eta^{j-1}\|^{2}\\[8.0pt] \quad-2j(j+1)\left\langle\displaystyle\frac{1}{j+1}(z^{0}-P_{c}(z^{j-1}))+P_{c}(z^{j-1})-z^{j-1},\displaystyle\frac{j}{j+1}\eta^{j-1}\right\rangle,\end{array}$ (3.8)
@@
$\begin{array}[]{l}\displaystyle\sum_{j=1}^{k}[A_{2}(j)+B_{2}(j)]=2(k+1)\langle z^{k}-P_{c}(z^{k}),z^{k}-z^{0}\rangle\\[6.0pt] ~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~\quad+2\displaystyle\sum_{j=1}^{k-1}(j+1)\langle z^{j}-P_{c}(z^{j}),z^{j}-P_{c}(z^{j})\rangle+2\|z^{0}-P_{c}(z^{0})\|^{2}.\end{array}$ (3.10)
@@
$D_{2}+E_{1}=-2k\left\langle z^{k}-z^{0},\eta^{k-1}\right\rangle-\displaystyle\sum_{j=0}^{k-1}2\left\langle z^{0}-z^{j},\frac{j+1}{j+2}\eta^{j}\right\rangle-\displaystyle\sum_{j=0}^{k-1}2j\left\langle z^{0}-z^{j},\frac{j+1}{j+2}\eta^{j}-\eta^{j-1}\right\rangle.$ (3.15)

### v2

@@ Abstract
This paper investigates a Halpern acceleration of the inexact proximal point method for solving maximal monotone inclusion problems in Hilbert spaces. The proposed Halpern inexact proximal point method (HiPPM) is shown to be globally convergent, and a unified framework is developed to analyze its worst-case convergence behavior. Under mild conditions on the inexactness tolerances, HiPPM achieves an $\mathcal{O}(1/k^{2})$ convergence rate in terms of the squared fixed-point residual. Moreover, under additional well-studied regularity conditions, the method attains a fast linear convergence rate. Building on this framework, we further extend the Halpern acceleration to the inexact augmented Lagrangian method for constrained convex optimization. In the spirit of Rockafellar’s classical results, the resulting accelerated inexact augmented Lagrangian method inherits the convergence rate and iteration complexity guarantees of HiPPM. Numerical experiments are provided to support the theoretical findings.
@@ 1 Introduction
where $T:\mathcal{H}\rightrightarrows\mathcal{H}$ is a maximally monotone operator. This problem provides a unifying framework for a broad range of models in optimization and equilibrium theory. In particular, many convex optimization problems, variational inequalities, and saddle-point formulations can be recast as finding the zeros of a monotone operator. The proximal point method (PPM) is a fundamental approach for solving monotone inclusion problems of the form (1.1). It not only enjoys strong global convergence guarantees for maximal monotone operators, but also serves as the theoretical foundation for various operator-splitting schemes. In particular, for convex optimization problems, both the augmented Lagrangian method (ALM) and the alternating direction method of multipliers (ADMM) can be interpreted as instances of the PPM when applied to their respective dual formulations.
@@
$P_{c}(z)=(I+cT)^{-1}(z),z\in{\cal H}.$
It follows from [18, Proposition 1] that $P_{c}(\cdot)$ is a nonexpansive mapping with
$0\in T(z)\mbox{ if and only if }z=P_{c}(z).$
Equivalently,
@@
In contrast to Nesterov-type acceleration, which was originally developed for convex optimization, Halpern’s fixed-point iteration [8] was introduced for approximating fixed points of nonexpansive operators. Recent studies have revealed a close connection between accelerated proximal algorithms and Halpern’s iteration. In particular, Contreras and Cominetti [5] established a precise relationship between Kim’s accelerated proximal point method [9] and Halpern’s scheme, suggesting that Halpern-type iterations can serve as a viable mechanism for accelerating the proximal point method. More recently, a Halpern-type accelerated proximal point method with a positive semidefinite preconditioner was studied in [19], this approach is shown to be equivalent to an accelerated preconditioned alternating direction method of multipliers and demonstrates strong numerical performance.
From a theoretical perspective, Lieder [11] analyzed Halpern’s iteration in Hilbert spaces and established an $\mathcal{O}(1/k^{2})$ convergence rate. Moreover, under a cocoercivity assumption, Tran-Dinh [20] showed an equivalence between Nesterov’s accelerated method and Halpern’s fixed-point iteration. Despite these advances, the existing connections are largely confined to exact algorithmic settings and do not directly extend to scenarios in which the proximal subproblems are solved inexactly, as is typically the case in large-scale applications. On the practical side, Halpern-type iterations have recently been employed to enhance GPU-based first-order methods, including preconditioned alternating direction method of multipliers, the Peaceman–Rachford method, and primal–dual hybrid gradient methods, for solving large-scale linear and quadratic programming problems; see, for example, [1, 13, 12, 19, 3, 4, 6].
In this paper, we study a Halpern-type acceleration of the inexact proximal point method (HiPPM) for solving the maximal monotone inclusion problem (1.1) in Hilbert spaces. The main contributions of this work can be summarized as follows:
@@
    As pointed out in [9], for strongly monotone operators, the standard proximal point method enjoys a linear convergence rate in terms of the fixed-point residual, whereas its Halpern-accelerated variant is not necessarily guaranteed to preserve such a rate. In this paper, we demonstrate that the HiPPM indeed attains linear convergence under some well-studied regularity conditions.
@@
The remainder of this paper is organized as follows. Section 2 introduces the framework of the Halpern accelerated inexact proximal point algorithm in Hilbert spaces and establishes its global convergence. Section 3 develops a general framework for analyzing the worst-case convergence rate of the proposed HiPPM via a direct algebraic proof, and further demonstrates its fast linear convergence under mild conditions. Section 4 derives the convergence rate of the Halpern accelerated inexact augmented Lagrangian method by exploiting its close connection with the Halpern accelerated inexact proximal point algorithm. Section 5 presents numerical experiments that validate the theoretical results. Finally, Section 6 concludes the paper.
@@
2 Halpern acceleration of the inexact PPM
Building upon Rockafellar’s inexact proximal point algorithm [17], we introduce an additional Halpern-type acceleration step to enhance its convergence behavior. The resulting accelerated scheme, referred to as HiPPM, can be formulated as follows.
@@ 2 Halpern acceleration of the inexact PPM
We recall the following result from [18, Proposition 2], which serves as a fundamental criterion for the existence of solutions to the inclusion problem $0\in T(z)$.
@@
For notational convenience, for $k=0,1,\ldots$, define
$\eta^{k}=\bar{z}^{k}-P_{c_{k}}(z^{k}).$
@@
$\|\bar{z}^{k}\|=\left\|\displaystyle\frac{k+2}{k+1}z^{k+1}-\displaystyle\frac{1}{k+1}z^{0}\right\|\leq 1.4s+\left\|\displaystyle\frac{1}{k+1}z^{0}\right\|<1.5s+0.5s=2s,$
@@
Then, the result follows by applying the same technique as in [18, Theorem 1]. For completeness, the details are provided below. Consider the set-valued mapping $T^{\prime}$ defined by
@@
Thus the sequence $\{z^{k}\}$ can be regarded as arising from the Halpern accelerated proximal point algorithm for operator $T^{\prime}$. As the effective domain ${\rm Dom}\,(T^{\prime})$ is bounded, so that $(T^{\prime})^{-1}(0)\neq\emptyset$ by Lemma 2.1. It follows from (2.6) that $(T^{\prime})^{-1}(0)\subseteq T^{-1}(0)$, this implies $T^{-1}(0)\neq\emptyset$. The proof is completed. ∎
@@
$\lim_{k\rightarrow\infty}\|Q_{c_{k}}(z^{k})\|=0,\quad\lim_{k\rightarrow\infty}\|\bar{z}^{k}-z^{k}\|=0.$
@@
$\|z^{k}\|<s\mbox{ and }\varepsilon_{k}<s,\quad\forall k.$
@@
$\|\bar{z}^{k}\|\leq 2s,\quad\|P_{c_{k}}(z^{k})\|\leq 3s,\quad\mbox{ when }k>N_{0}.$
@@
$\begin{array}[]{l}\|Q_{c_{k}}(z^{k})\|^{2}-\|z^{k}-z^{0}\|^{2}+\|z^{k+1}-z^{0}\|^{2}\\[8.0pt] \leq\|z^{k+1}-z^{0}\|^{2}-\|P_{c_{k}}(z^{k})-z^{0}\|^{2}\\[8.0pt] =\left(\displaystyle\frac{k+1}{k+2}\right)^{2}\|z^{0}-\bar{z}^{k}\|^{2}-\|(P_{c_{k}}(z^{k})-\bar{z}^{k})+(\bar{z}^{k}-z^{0})\|^{2}\\[8.0pt] \leq-\|(P_{c_{k}}(z^{k})-\bar{z}^{k})\|^{2}-2\langle(P_{c_{k}}(z^{k})-\bar{z}^{k}),(\bar{z}^{k}-z^{0})\rangle\\[8.0pt] \leq 2\|P_{c_{k}}(z^{k})-\bar{z}^{k}\|\|\bar{z}^{k}-z^{0}\|\leq 2(\|z^{0}\|+2s)\varepsilon_{k}.\end{array}$
@@
$\begin{array}[]{l}\|Q_{c_{k}}(z^{k})\|^{2}-\|z^{k}-\bar{z}\|^{2}+\|z^{k+1}-\bar{z}\|^{2}\\[8.0pt] \leq\|z^{k+1}-\bar{z}\|^{2}-\|P_{c_{k}}(z^{k})-\bar{z}\|^{2}\\[8.0pt] =\left\|\displaystyle\frac{1}{k+2}(z^{0}-\bar{z})+\displaystyle\frac{k+1}{k+2}(\bar{z}^{k}-\bar{z})\right\|^{2}-\|P_{c_{k}}(z^{k})-\bar{z}^{k}+\bar{z}^{k}-\bar{z}\|^{2}\\[8.0pt] \leq\displaystyle\frac{1}{(k+2)^{2}}\|(z^{0}-\bar{z})\|^{2}-2\langle P_{c_{k}}(z^{k})-\bar{z}^{k},\bar{z}^{k}-\bar{z}\rangle-\|P_{c_{k}}(z^{k})-\bar{z}^{k}\|^{2}\\[8.0pt] \leq\displaystyle\frac{1}{(k+2)^{2}}\|(z^{0}-\bar{z})\|^{2}+2(2s+\|\bar{z}\|)\varepsilon_{k}.\end{array}$
@@
Arguing as in the proof of [18, Theorem 1], one can show that the sequence $\{z^{k}\}$ admits at most one weak cluster point. Hence, the entire sequence $\{z^{k}\}$ converges weakly to some $z^{\infty}$. ∎
@@ 3 The rate of convergence
$\begin{array}[]{ll}\Delta_{k}&=\displaystyle\frac{4k}{(k+1)^{2}}\left\langle z^{k}-z^{0},\eta^{k-1}\right\rangle+\displaystyle\frac{4}{(k+1)^{2}}\displaystyle\sum_{j=0}^{k-1}\left\langle z^{0}-z^{j},\frac{j+1}{j+2}\eta^{j}\right\rangle\\[6.0pt] &\quad+\displaystyle\frac{4}{(k+1)^{2}}\displaystyle\sum_{j=0}^{k-1}j\left\langle z^{0}-z^{j},\frac{j+1}{j+2}\eta^{j}-\eta^{j-1}\right\rangle+\displaystyle\frac{4k}{k+1}\left\langle z^{k}-P_{c}(z^{k}),\eta^{k-1}\right\rangle\\[6.0pt] &\quad-\displaystyle\frac{4}{(k+1)^{2}}\displaystyle\sum_{j=0}^{k-1}\displaystyle\frac{(j+1)^{3}}{j+2}\left\langle P_{c}(z^{j})-z^{j},\eta^{j-1}-\eta^{j}\right\rangle\\[6.0pt] &\quad+\displaystyle\frac{4}{(k+1)^{2}}\displaystyle\sum_{j=0}^{k-1}\displaystyle\frac{j+1}{j+2}\left\langle P_{c}(z^{j})-z^{j},\eta^{j-1}\right\rangle-\displaystyle\frac{4}{(k+1)^{2}}\displaystyle\sum_{j=1}^{k}\displaystyle\frac{j^{2}}{j+1}\|\eta^{j-1}\|^{2}.\end{array}$ (3.14)
@@
$\begin{array}[]{l}j(j+1)\|P_{c}(z^{j})-P_{c}(z^{j-1})\|^{2}\\[4.0pt] =j(j+1)\left\|z^{j}-P_{c}(z^{j})+\displaystyle\frac{1}{j}[z^{j}-z^{0}]-\eta^{j-1}\right\|^{2}\\[4.0pt] =j(j+1)\left[\left\|z^{j}-P_{c}(z^{j})+\displaystyle\frac{1}{j}[z^{j}-z^{0}]\right\|^{2}+\|\eta^{j-1}\|^{2}-2\langle z^{j}-P_{c}(z^{j})+\displaystyle\frac{1}{j}[z^{j}-z^{0}],\eta^{j-1}\rangle\right]\\[4.0pt] =j(j+1)\|z^{j}-P_{c}(z^{j})\|^{2}+2(j+1)\langle z^{j}-P_{c}(z^{j}),z^{j}-z^{0}\rangle+\displaystyle\frac{j+1}{j}\|z^{j}-z^{0}\|^{2}\\[8.0pt] \quad-2j(j+1)\langle z^{j}-P_{c}(z^{j})+\displaystyle\frac{1}{j}[z^{j}-z^{0}],\eta^{j-1}\rangle+j(j+1)\|\eta^{j-1}\|^{2}\\[6.0pt] =A_{1}(j)+A_{2}(j)+A_{3}(j)-2j(j+1)\langle z^{j}-P_{c}(z^{j})+\displaystyle\frac{1}{j}[z^{j}-z^{0}],\eta^{j-1}\rangle+j(j+1)\|\eta^{j-1}\|^{2},\end{array}$
@@
$\begin{array}[]{ll}A_{1}(j)=j(j+1)\|z^{j}-P_{c}(z^{j})\|^{2},\,\,A_{2}(j)=2(j+1)\langle z^{j}-P_{c}(z^{j}),z^{j}-z^{0}\rangle,\\[4.0pt] \hbox{and}A_{3}(j)=\displaystyle\frac{j+1}{j}\|z^{j}-z^{0}\|^{2}.\end{array}$
@@
$\begin{array}[]{l}-j(j+1)\|z^{j}-z^{j-1}\|^{2}\\[4.0pt] =-j(j+1)\left\|\displaystyle\frac{1}{j+1}(z^{0}-P_{c}(z^{j-1}))+P_{c}(z^{j-1})+\displaystyle\frac{j}{j+1}\eta^{j-1}-z^{j-1}\right\|^{2}\\[8.0pt] =-2j\langle z^{0}-P_{c}(z^{j-1}),P_{c}(z^{j-1})-z^{j-1}\rangle-j(j+1)\|P_{c}(z^{j-1})-z^{j-1}\|^{2}\\[8.0pt] \quad-\displaystyle\frac{j}{j+1}\|z^{0}-P_{c}(z^{j-1})\|^{2}-2j(j+1)\left\langle\displaystyle\frac{1}{j+1}(z^{0}-P_{c}(z^{j-1}))+P_{c}(z^{j-1})-z^{j-1},\displaystyle\frac{j}{j+1}\eta^{j-1}\right\rangle\\[8.0pt] \quad-j(j+1)\left(\displaystyle\frac{j}{j+1}\right)^{2}\|\eta^{j-1}\|^{2}\\[8.0pt] =B_{1}(j)+B_{2}(j)+B_{3}(j)-j(j+1)\left(\displaystyle\frac{j}{j+1}\right)^{2}\|\eta^{j-1}\|^{2}\\[8.0pt] \quad-2j(j+1)\left\langle\displaystyle\frac{1}{j+1}(z^{0}-P_{c}(z^{j-1}))+P_{c}(z^{j-1})-z^{j-1},\displaystyle\frac{j}{j+1}\eta^{j-1}\right\rangle,\end{array}$
@@
$\begin{array}[]{l}\displaystyle\sum_{j=1}^{k}[A_{2}(j)+B_{2}(j)]=2(k+1)\langle z^{k}-P_{c}(z^{k}),z^{k}-z^{0}\rangle\\[6.0pt] ~~~~~~~~~~~~~~~~~~~~~~~~~\quad+2\displaystyle\sum_{j=1}^{k-1}(j+1)\langle z^{j}-P_{c}(z^{j}),z^{j}-P_{c}(z^{j})\rangle+2\|z^{0}-P_{c}(z^{0})\|^{2}.\end{array}$ (3.20)
@@
$D_{2}+E_{1}=-2k\left\langle z^{k}-z^{0},\eta^{k-1}\right\rangle-\displaystyle\sum_{j=0}^{k-1}2\left\langle z^{0}-z^{j},\frac{j+1}{j+2}\eta^{j}\right\rangle-\displaystyle\sum_{j=0}^{k-1}2j\left\langle z^{0}-z^{j},\frac{j+1}{j+2}\eta^{j}-\eta^{j-1}\right\rangle.$